Please solve.
4(x+2)=

Answer by jim_thompson5910(35256)

Use the distributive property to get


4(x+2) = 4(x)+4(2) = 4x+8


So 4(x+2) = 4x+8


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim